Some quarters and nickels have a total value of $6.05. Find the number of each if there are two times as many quarters as nickels.
Answer by ikleyn(52809)

.

x nickels

2x quarters


The money equation is

    5x + 25(2x) = 605   cents

    5x +  50x   = 605

       55x      = 605

         x      = 605/55 = 11.


ANSWER.  11 nickels and 22 quarters.


CHECK.  11*5 + 22*25 = 605  cents,  total.   ! correct !

Solved.